A computer repair bill was $225. This included $75 for parts and $50 for each hour of labor. Find the number of hours of labor.
Found 2 solutions by Paul, chripchrip:
Answer by Paul(988)

75+50x=225
50x=150
x=3

Hence the number of hours was 3
Paul.

Answer by chripchrip(6)

let x be the labour hour
75+50x=225
50x=225-75
50x=150
x=3
the nos of labour hour is 3